OpenStudy (anonymous):

If f(x) = log2 (x + 4), what is f^(−1)(3)?

OpenStudy (anonymous):

Do you mean \[f(x) = \log_{2} (x+4)\] \[f ^{-1}(x) = ?\]

OpenStudy (anonymous):

Correction... \[f ^{-1}(3) = ?\]

OpenStudy (jdoe0001):

\(\bf f(x) =y= log_2(x + 4)\qquad inverse\implies x= log_2(y + 4)\\ \quad \\ \quad \\ \textit{log cancellation rule of }a^{log_ax}=x\qquad thus\\ \quad \\ x= log_2(y + 4)\implies 2^x=2^{log_2(y + 4)}\implies 2^x=y+4 \)

OpenStudy (jdoe0001):

solve for "y" and plug in "3" for \(\bf f^{-1}\)

OpenStudy (anonymous):

yes thats what i meant! thank you

OpenStudy (anonymous):

where would f^-1 be though

OpenStudy (jdoe0001):

?

OpenStudy (jdoe0001):

well.. what did you get for the \(\bf f^{-1}\quad ?\)

OpenStudy (anonymous):

so when you solve for "y" you get 2^x - 4 = y

OpenStudy (isaiah.feynman):

Yes!

OpenStudy (anonymous):

f^-1 would be 9? cause 3 times three?

OpenStudy (jdoe0001):

f(9) just means , make "x" equals to 9 for all intances of "x" the same is true for \(\bf f^{-1}(9)\)

OpenStudy (jdoe0001):

\(\bf 2^x-4=y\qquad \qquad f^{-1}(\color{red}{9})=2^{\color{red}{9}}-4\)

OpenStudy (anonymous):

so how would i get the answer from that?

OpenStudy (isaiah.feynman):

The answer is already looking at you, what is 2 to the 9? punch your calculator whatever you get subtract 4 from it and thats your answer!